OpenStudy (anonymous):

Mrs.collins is organizing a school trip. 120 students are going on the trip. the ratio of the number of the staff to the number of students must be 1:15. Mrs Collins book three coaches for the trip. each coach has 42 seats. Are there enough seats for all the students and the staff?. pls show the working

Miracrown (miracrown):

so if there are 120 students and the ratio of staff to students is 1:15 then how to solve the number of staff?

Miracrown (miracrown):

any idea? @s4

OpenStudy (arindameducationusc):

@Miracrown Should I try?

OpenStudy (anonymous):

nope no idea

Miracrown (miracrown):

Remember proportion?

Miracrown (miracrown):

so in this case number of staff is proportion to number of student by a ratio of 1:15 so 1 is proportion to 15 as if what is proportion to what? remember number of staff i let it to be x and number of student is given to be 120 so 1:15

OpenStudy (arindameducationusc):

x/120=1/15, we get x=8 (teachers) now, 42*3coaches=126 but student+teachers, i.e, 120+8=128

OpenStudy (anonymous):

ohh

Miracrown (miracrown):

it is essentially saying if there are 15 student then 1 staff is taking care of them, now there are 120 students, how many staff is needed to take care of those 120 students?

OpenStudy (anonymous):

I get it now, thank you :D

Miracrown (miracrown):

So what is it? :D

OpenStudy (anonymous):

So this is wat I did 1+15=16 120/16=7.5=8 1x8=8(for staff) + 15x8=120(for students) =128 now for coaches 42x3=126 so its not enough they need 2 more seats

OpenStudy (anonymous):

is it right ?